OpenStudy (anonymous):

how do i isolate c with the question a=b(1/c-1/d)

OpenStudy (anonymous):

Looks like a Maths question..:/

OpenStudy (anonymous):

If it's Maths question..than it would be better ..that you ask this in Maths section!

OpenStudy (anonymous):

http://openstudy.com/study#/groups/Mathematics

OpenStudy (alienium):

a = b ( 1/c-1/d) a = b ( 1/c-1 * d/1) a = b * d/c-1 c-1 = b * d/a c = (b * d/a) +1